# Exact WKB in all sectors I: Potentials with degenerate saddles

### v1

@@ 5.2 The duality
Let us now discuss the duality between the bottom of wells and the barrier top of the symmetric double-well potential. As we stated above, the dual theory is the unstable quartic oscillator. Therefore, the expressions for both theories are not the same and the transformations between them are non-trivial contrary to the periodic potential case. For example, we express the series expansions of $\mathcal{F}$ and $\mathcal{G}$ around $u\sim 0$ and $u\sim u_{\mathrm{T}}$ first few orders in $g$ as
$\mathcal{F}^{u\sim 0}(\tilde{u},g)$ $=\tilde{u}+g\left(\frac{1}{8}+\frac{3\tilde{u}^{2}}{2}\right)+g^{2}\left(\frac{25\tilde{u}}{16}+\frac{35\tilde{u}^{3}}{4}\right)$
$\;\;+g^{3}\left(\frac{175}{256}+\frac{735\tilde{u}^{2}}{32}+\frac{1155\tilde{u}^{4}}{16}\right)+\dots,$ (179)
$\mathcal{G}^{u\sim 0}(\tilde{u},g)$ $=\frac{1}{3g}+g\left(\frac{19}{48}+\frac{17\tilde{u}^{2}}{4}\right)+g^{2}\left(\frac{187\tilde{u}}{32}+\frac{227\tilde{u}^{3}}{8}\right)$
$\;\;+g^{3}\left(\frac{28829}{9216}+\frac{34121\tilde{u}^{2}}{384}+\frac{47431\tilde{u}^{4}}{192}\right)+\dots,$ (180)
@@
$\mathcal{F}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)$ $=i\sqrt{2}\tilde{u}+g\left(\frac{3i}{8\sqrt{2}}-\frac{3i\tilde{u}^{2}}{\sqrt{2}}\right)+g^{2}\left(-\frac{85i\tilde{u}}{16\sqrt{2}}+\frac{35i\tilde{u}^{3}}{2\sqrt{2}}\right)$
$\;\;+g^{3}\left(-\frac{1995i}{512\sqrt{2}}+\frac{2625i\tilde{u}^{2}}{32\sqrt{2}}-\frac{1155i\tilde{u}^{4}}{8\sqrt{2}}\right)+\dots,$ (181)
$\mathcal{G}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)$ $=-\frac{i}{3\sqrt{2}g}+g\left(\frac{67i}{96\sqrt{2}}-\frac{17i\tilde{u}^{2}}{4\sqrt{2}}\right)+g^{2}\left(-\frac{671i\tilde{u}}{64\sqrt{2}}+\frac{227i\tilde{u}^{3}}{8\sqrt{2}}\right)$
$\;\;+g^{3}\left(-\frac{372101i}{36864\sqrt{2}}+\frac{125333i\tilde{u}^{2}}{768\sqrt{2}}-\frac{47431i\tilde{u}^{4}}{192\sqrt{2}}\right)+\dots.$ (182)
Then, we observe that while $\mathcal{F}^{u\sim 0}$ and $\mathcal{G}^{u\sim 0}$ obey the P-NP relation around $u\sim 0$ which is given in (159), while $\mathcal{F}^{u\sim u_{\mathrm{T}}}$ and $\mathcal{G}^{u\sim u_{\mathrm{T}}}$ satisfy a slightly different P-NP relation:
$\mathcal{G}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)=-\int\frac{{\mathrm{d}}g}{g}\,\left[\sqrt{2}\,\frac{S_{\mathrm{I}}^{u\sim u_{\mathrm{T}}}}{ig}\frac{\partial\tilde{u}\left(\mathcal{F}^{u\sim u_{\mathrm{T}}},g\right)}{\partial\mathcal{F}^{u\sim u_{\mathrm{T}}}}+\frac{\mathcal{F}^{u\sim u_{\mathrm{T}}}}{\sqrt{2}}\right]\,,$ (183)
@@
$S^{u\sim u_{\mathrm{T}}}=\int_{0}^{1}{\mathrm{d}}x\,\sqrt{2\left(V(x)-1\right)}=\frac{i}{3\sqrt{2}}\,.$ (184)
@@
$S^{u\sim u_{\mathrm{T}}}=\frac{a_{0}(1)}{2}=\lim_{g\rightarrow 0}\left[g\,\mathcal{G}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)\right]\,.$
@@
$\mathcal{G}^{\mathrm{D}}(\tilde{u},g)=-\int\frac{{\mathrm{d}}g}{g}\,\left[\sqrt{2}\frac{S_{\mathrm{I}}^{\mathrm{D}}}{g}\frac{\partial\tilde{u}\left(\mathcal{F}^{\mathrm{D}},g\right)}{\partial\mathcal{F}^{\mathrm{D}}}+\frac{\mathcal{F}^{\mathrm{D}}}{\sqrt{2}}\right]\,,$ (186)
@@
is equal to the half bounce action of the unstable quartic oscillator. The P-NP relation in (186) is for the unstable quartic oscillator and it is indeed related to (159) via the transformations given in (114) for $\k{_}\mathrm{DW}=2_$, i.e.
$\mathcal{F}\rightarrow\frac{\mathcal{F}}{\sqrt{2}}\,,\qquad\mathcal{G}\rightarrow\sqrt{2}\mathcal{G}^{\mathrm{D}}\,,\qquad S_{\mathrm{I}}\rightarrow\sqrt{2}S_{\mathrm{I}}^{\mathrm{D}}\,.$ (188)
Note that as we discussed in Section 3 $\k{_}{\mathrm{DW}}=2_$ is the ratio of the classical (dual) actions of both theories. Equivalently, in the Weber-type EWKB framework, it can also be deduced from the relations
$\mathcal{F}_{0}\Big{(}\frac{u}{g}\Big{)}=\sqrt{2}\mathcal{F}_{0}\Big{(}\frac{u}{g}\Big{)}\,,\qquad\mathcal{G}_{0}^{\mathrm{D}}\Big{(}\frac{u}{g}\Big{)}=\frac{\mathcal{G}_{0}\Big{(}\frac{u}{g}\Big{)}}{\sqrt{2}}\,.$ (189)
@@ 6 Discussion and outlook
  • Generalization of the EWKB Method: In Section 2, we discuss the generalization of the EWKB method to incorporate all sectors and saddle points for locally harmonic potentials. We focused on two distinct complementary approaches: the Airy-type and Weber-type EWKB methods.
    In Section 2.1, introducing the complexification of the parameter $u$, we extended the application realm of the Airy-type approach, which was historically limited to below-barrier sectors of potentials, to include regions above the barrier. This extension provides a systematic way to track the quantization conditions through analytical continuation across different sectors for generic locally harmonic potentials.
    In Section 2.2, we examined the Weber-type EWKB, which provides exact estimations of the quantum action around saddle points. Unlike the Airy-type method, which is more qualitative, the Weber-type method allows for precise quantitative estimates. Along with identifying local and non-local components of quantum actions, we extended their analysis to all saddle points which presents a generalization of Zinn-Justin and Jentschura’s results Zinn-Justin:2004qzw ; Zinn-Justin:2004vcw ; Jentschura:2010zza ; Jentschura:2011zza . Notably, we identified an exchange between perturbative and non-perturbative behaviors near minima and maxima, which is controlled by $S$-duality –a powerful symmetry that links weak and strong coupling regimes in quantum systems.
@@
    The transition between different potential sectors revealed the role of $S$-duality in maintaining consistency between perturbative and non-perturbative contributions. This duality effectively swaps the behavior around minima and maxima, offering insights into the system’s symmetry structure. Importantly, we found that the duality can be understood in terms of Weber-type EWKB elements, namely the functions $\mathcal{F}$ and $\mathcal{G}$, further extending its utility in analyzing complex quantum mechanical systems.
@@
  • Resurgence and $S$-Duality:  Another central theme in our study is the resurgence theory which exhibits the deep connection between perturbative and non-perturbative sectors. Through carefully tracking of quantum actions across different sectors of potentials and using the Stokes automorphism and the median resummation along with appropriate analytic continuations, we showed that the resurgence structure is preserved in all sectors. This preservation is particularly important for systems where classical approximations break down, such as near turning points or in the presence of degenerate minima.
@@
    By extending the Weber-type method to handle all saddle points, we were also able to generalize the acknowledged results around multiple degenerate minima to all type of degenerate saddle points. This generalization is particularly relevant for understanding the role of non-perturbative effects in complex quantum systems. The exact estimates of quantum actions we obtained could be applied to a variety of systems, ranging from simple harmonic oscillators to more exotic potentials with rich symmetry structures.
@@ Acknowledgements
This work of T. M. is supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research (KAKENHI) Grant Numbers 23K03425 and 22H05118. C.P. thanks the hospitality of Kindai University where the initial steps of the project took place. The beginning of this work is also owed to the discussion in the workshop ‘Invitation to Recursion, Resurgence and Combinatorics’ at Okinawa Institute of Science and Technology Graduate University (OIST).
@@ Appendix B Notation and other definitions
In this appendix, we provide a list of variables which we use in the main text to clarify the notation and definitions.
@@
$s$: Ricatti function $\Omega$: Ricatti function $S$: Ricatti function
@@
$\theta_{\lambda}$: Phase for the parameter $\lambda$ $\Psi_{\mathrm{L}}$: Local wave function $\Psi_{\mathrm{G}}$: Global wave function

### v2

@@ 5.2 The duality
Let us now discuss the duality between the bottom of wells and the barrier top of the symmetric double-well potential. As we stated above, the dual theory is the unstable quartic oscillator. The perturbative and non-perturbative quantities for these theories are not the same and the transformations between them are non-trivial contrary to the periodic potential case. For example, we get the series expansions of $\mathcal{F}$ and $\mathcal{G}$ around $u\sim 0$ and $u\sim u_{\mathrm{T}}$ first few orders in $g$ as
$\mathcal{F}^{u\sim 0}(\tilde{u},g)$ $=\tilde{u}+g\left(\frac{1}{8}+\frac{3\tilde{u}^{2}}{2}\right)+g^{2}\left(\frac{25\tilde{u}}{16}+\frac{35\tilde{u}^{3}}{4}\right)+g^{3}\left(\frac{175}{256}+\frac{735\tilde{u}^{2}}{32}+\frac{1155\tilde{u}^{4}}{16}\right)$
$\;\;+g^{4}\left(\frac{31185\tilde{u}}{1024}+\frac{45045\tilde{u}^{3}}{128}+\frac{45045\tilde{u}^{5}}{64}\right)+\dots,$ (180)
$\mathcal{G}^{u\sim 0}(\tilde{u},g)$ $=\frac{2}{3g}+g\left(\frac{19}{24}+\frac{17\tilde{u}^{2}}{2}\right)+g^{2}\left(\frac{187\tilde{u}}{16}+\frac{227\tilde{u}^{3}}{4}\right)+g^{3}\left(\frac{28829}{4608}+\frac{34121\tilde{u}^{2}}{192}+\frac{47431\tilde{u}^{4}}{96}\right)$
$\;\;+g^{4}\left(\frac{842909\tilde{u}}{3072}+\frac{264725\tilde{u}^{3}}{96}+\frac{317629\tilde{u}^{5}}{64}\right)+\dots,$ (181)
@@
$\mathcal{F}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)$ $=i\sqrt{2}\tilde{u}+g\left(\frac{3i}{8\sqrt{2}}-\frac{3i\tilde{u}^{2}}{\sqrt{2}}\right)+g^{2}\left(-\frac{85i\tilde{u}}{16\sqrt{2}}+\frac{35i\tilde{u}^{3}}{2\sqrt{2}}\right)+g^{3}\left(-\frac{1995i}{512\sqrt{2}}+\frac{2625i\tilde{u}^{2}}{32\sqrt{2}}-\frac{1155i\tilde{u}^{4}}{8\sqrt{2}}\right)$
$\;\;+g^{4}\left(\frac{400785i\tilde{u}}{2048\sqrt{2}}-\frac{165165i\tilde{u}^{3}}{128\sqrt{2}}+\frac{45045i\tilde{u}^{5}}{32\sqrt{2}}\right)+\dots,$ (182)
$\mathcal{G}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)$ $=-\frac{i\sqrt{2}}{3g}+g\left(\frac{67i}{48\sqrt{2}}-\frac{17i\tilde{u}^{2}}{2\sqrt{2}}\right)+g^{2}\left(-\frac{671i\tilde{u}}{32\sqrt{2}}+\frac{227i\tilde{u}^{3}}{4\sqrt{2}}\right)+$
$\;\;+g^{3}\left(-\frac{372101i}{18432\sqrt{2}}+\frac{125333i\tilde{u}^{2}}{384\sqrt{2}}-\frac{47431i\tilde{u}^{4}}{96\sqrt{2}}\right)$
$\;\;+g^{4}\left(\frac{3839943i\tilde{u}}{4096\sqrt{2}}-\frac{82315i\tilde{u}^{3}}{16\sqrt{2}}+\frac{317629i\tilde{u}^{5}}{64\sqrt{2}}\right)+\dots\;.$ (183)
This recovers the expressions in DP1 and adds a couple of more orders. Note that for our purpose, the series of $\mathcal{F}$ and $\mathcal{G}$ are important due to their relationship via the P-NP relation and establishing the $S$-transformation for the latter. In that sense, we observe that while $\mathcal{F}^{u\sim 0}$ and $\mathcal{G}^{u\sim 0}$ obey the P-NP relation around $u\sim 0$ which is given in (160), $\mathcal{F}^{u\sim u_{\mathrm{T}}}$ and $\mathcal{G}^{u\sim u_{\mathrm{T}}}$ satisfy a slightly different P-NP relation:
$\mathcal{G}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)=-2\int\frac{{\mathrm{d}}g}{g}\,\left[\sqrt{2}\,\frac{S_{\mathrm{I}}^{u\sim u_{\mathrm{T}}}}{ig}\frac{\partial\tilde{u}\left(\mathcal{F}^{u\sim u_{\mathrm{T}}},g\right)}{\partial\mathcal{F}^{u\sim u_{\mathrm{T}}}}+\frac{\mathcal{F}^{u\sim u_{\mathrm{T}}}}{\sqrt{2}}\right]\,,$ (184)
@@
$S_{\mathrm{I}}^{u\sim u_{\mathrm{T}}}=\int_{0}^{1}{\mathrm{d}}x\,\sqrt{2\left(V(x)-u_{\mathrm{T}}\right)}=\frac{i}{3\sqrt{2}}\,.$ (185)
@@
$S_{\mathrm{I}}^{u\sim u_{\mathrm{T}}}=\frac{a_{0}(u_{\mathrm{T}})}{2}=\lim_{g\rightarrow 0}\left[g\,\mathcal{G}^{u\sim u_{\mathrm{T}}}(\tilde{u},g)\right]\,.$
@@
$\mathcal{G}^{\mathrm{D}}(\tilde{u},g)=-2\int\frac{{\mathrm{d}}g}{g}\,\left[\sqrt{2}\frac{S_{\mathrm{I}}^{\mathrm{D}}}{g}\frac{\partial\tilde{u}\left(\mathcal{F}^{\mathrm{D}},g\right)}{\partial\mathcal{F}^{\mathrm{D}}}+\frac{\mathcal{F}^{\mathrm{D}}}{\sqrt{2}}\right]\,,$ (187)
@@
is equal to the half bounce action of the unstable quartic oscillator. The P-NP relation in (187) is for the unstable quartic oscillator and it is indeed related to (160) via the transformations given in (114) for $\k{_}2=2_$, i.e.
$\mathcal{F}\rightarrow\frac{\mathcal{F}^{\mathrm{D}}}{\sqrt{2}}\,,\qquad\mathcal{G}\rightarrow\sqrt{2}\mathcal{G}^{\mathrm{D}}\,,\qquad S_{\mathrm{I}}\rightarrow\sqrt{2}S_{\mathrm{I}}^{\mathrm{D}}\,.$ (189)
Note that as we discussed in Section 3, while $\k{_}2=2_$ is the ratio of the classical (dual) actions of both theories but it is not the case for higher order terms. Equivalently, in the Weber-type EWKB framework, it can also be deduced from the relations
$\mathcal{F}_{0}^{\mathrm{D}}\Big{(}\frac{u}{g}\Big{)}=\sqrt{2}\mathcal{F}_{0}\Big{(}\frac{u}{g}\Big{)}\,,\qquad\mathcal{G}_{0}^{\mathrm{D}}\Big{(}\frac{u}{g}\Big{)}=\frac{\mathcal{G}_{0}\Big{(}\frac{u}{g}\Big{)}}{\sqrt{2}}\,.$ (190)
@@ 6 Discussion and outlook
  • Generalization of the EWKB Method: In Section 2, we discussed the generalization of the EWKB method to incorporate all sectors and saddle points for locally harmonic potentials. We focused on two distinct complementary approaches: the Airy-type and Weber-type EWKB methods.
    In Section 2.1, introducing the complexification of the parameter $u$, we extended the application realm of the Airy-type approach, which was historically limited to below-barrier sectors of potentials, to include regions above the barrier. This extension provides a systematic way to track the quantization conditions through analytic continuation across different sectors for generic locally harmonic potentials.
    In Section 2.2, we examined the Weber-type EWKB, which provides exact estimations of the quantum actions around saddle points. Unlike the Airy-type method, which is more qualitative, the Weber-type method allows for precise quantitative estimates. Along with identifying local and non-local components of quantum actions, we extended their analysis to all saddle points which presents a generalization of Zinn-Justin and Jentschura’s results Zinn-Justin:2004qzw ; Zinn-Justin:2004vcw ; Jentschura:2010zza ; Jentschura:2011zza . This generalization is particularly relevant for understanding the role of non-perturbative effects in complex quantum systems. The exact estimates of quantum actions we obtained could be applied to a variety of systems, ranging from simple harmonic oscillators to more exotic potentials with rich symmetry structures. Notably, we also identified an exchange between perturbative and non-perturbative behaviors near minima and maxima, which is controlled by $S$-duality –a powerful symmetry that links weak and strong coupling regimes in quantum systems.
@@
    The transition between different potential sectors revealed the role of $S$-duality in maintaining consistency between perturbative and non-perturbative contributions. This duality effectively swaps the behavior around minima and maxima, offering insights into the system’s symmetry structure. Importantly, we found that the duality can be understood in terms of the Weber-type EWKB elements, namely the functions $\mathcal{F}$ and $\mathcal{G}$, further extending its utility in analyzing complex quantum mechanical systems.
@@
  • Resurgence and $S$-Duality:  Another central theme in our study is the resurgence theory which exhibits the deep connection between perturbative and non-perturbative sectors. Through carefully tracking of quantum actions across different sectors of potentials and using the Stokes automorphism and the median resummation along with appropriate analytic continuations, we showed that the resurgence structure is preserved in all sectors.
@@ Acknowledgements
This work of T. M. is supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research (KAKENHI) Grant Numbers 23K03425 and 22H05118. C.P. thanks for the hospitality of Kindai University where the initial steps of the project took place. The beginning of this work is also owed to the discussion in the workshop ‘Invitation to Recursion, Resurgence and Combinatorics’ at Okinawa Institute of Science and Technology Graduate University (OIST). We also thank Alireza Behtash for his comments on the first manuscript.
@@ Appendix B Notation and other definitions
A list of variables which we use in the main text to clarify the notation and definitions.
@@
$D^{\mathrm{med}}=0$: Median quantization condition $\tau$: Modular Parameter
$\mathcal{F}(\tilde{u},g)$: Perturbative data in P-NP relation $\mathcal{G}(\tilde{u},g)$: Nonperturbative data in P-NP relation
@@
$s$: Riccati function $\Omega$: Riccati function $S$: Riccati function
@@
$\psi$: Wave function $\Psi_{\mathrm{L}}$: Local wave function $\Psi_{\mathrm{G}}$: Global wave function